OpenStudy (anonymous):

The volume of a cylinder is 24 cubic feet. What is the volume of a cone that fits exactly inside the cylinder? Provide an explanation and proof for your answer to receive full credit.

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

The volume of a cylinder is 24 cubic feet. What is the volume of a cone that fits exactly inside the cylinder? Provide an explanation and proof for your answer to receive full credit.

OpenStudy (anonymous):

This what I have so far am I doing it right??

OpenStudy (turingtest):

firstly, please post the actual question in the box so that people know if they can help you without clicking your question.

OpenStudy (anonymous):

volume of a cylinder = pi*r^2*h volume of cone = 1/3*pi*r^2*h same radius and height means same r and h given volume of cylinder = 24 so volume of cone = 1/3*24 = 8

OpenStudy (turingtest):

looks good

OpenStudy (turingtest):

i would have added one line to make it more formal: volume of a cylinder = pi*r^2*h volume of cone = 1/3*pi*r^2*h same radius and height means same r and h given volume of cylinder = 24 so volume of cone = 1/3*(volume of cylinder) = 24/3 = 8 ^^^^^^^^

OpenStudy (anonymous):

ok good can you help me explain it?

OpenStudy (turingtest):

i think that's a pretty good explanation... you want it in a more 'Englishy' format?

OpenStudy (anonymous):

yea thats what I want

OpenStudy (turingtest):

i guess something like... the formula for the volume of a cylinder depends only on height and radius, and differs from that for a cone by a factor of 1/3. hence if the height and radius are the same between them, the volume of the cone is 1/3 that of the cylinder

OpenStudy (turingtest):

i like your explanation better, but oh well

OpenStudy (anonymous):

Thanks I do to :)